OpenStudy (jamanorthi):

Solve 3x+12= 37-2x. help please!!!!

OpenStudy (anonymous):

x= 5

OpenStudy (jamanorthi):

but how??

OpenStudy (jamanorthi):

@heavenfun13 i need to show work

OpenStudy (anonymous):

3x+12=37-2x add 2x to both sides 5x+12=37 subtract 12 from both sides 5x=25 divide by 5 x=5

OpenStudy (jamanorthi):

thank you @juliannescriv

OpenStudy (anonymous):

mhm :)